OpenStudy (anonymous):

A gardener planted some flowers and kept a record of their height over the first 14 days of growth. Here is the chart: Time(days) 0 3 7 10 14 Height (cm) 4.2 6.3 9.1 11.2 14 Write an explicit rule for t(n) that gives the height of the flower after n days.

OpenStudy (allank):

I'm not sure but I think: t(n)=0.7n+4.2

OpenStudy (allank):

I've checked all of them. And now I'm sure that's the correct function.

OpenStudy (allank):

Do you know how to get the function? @maine2

OpenStudy (anonymous):

no

OpenStudy (allank):

First, have a look at the height at 0 days. It shows that a flower has a height of 4.2cm, with no time having elapsed. That gives us the constant of the function. t(n)=........+4.2 You comfortable with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would that be the rule

OpenStudy (anonymous):

t(n)=0.7n+4.2

OpenStudy (allank):

Yes

OpenStudy (anonymous):

ok

OpenStudy (allank):

You get the 0.7n this way: We write the function like t(n)= xn+4.2 and then plug in a height and its corresponding time. e.g. 3 days, 6.3cm 6.3=x*3+4.2 2.1=x*3 x=0.7 Plug in 0.7 into the function t(n)=0.7n+4.2 and that's the rule.

OpenStudy (anonymous):

ok i got it

OpenStudy (allank):

You're welcome :)